OpenStudy (anonymous):

Points V, W, X, Y, and Z are graphed on a number line. The coordinate of point V is –8. What points are at a distance of 5 units from point V? Choose all answers that are correct. A. W = –13 B. X = –8 C. Y = –3 D. Z = 3

OpenStudy (anonymous):

@YoungStudier

OpenStudy (texaschic101):

ok....your at -8 on a number line...and you need to find the points that are 5 units away...going in both directions. -8 + 5 = - 3 -8 + (-5) = -8 - 5 = - 13 so your points are -3 and - 13....Y and W